Suppose that on earth you can throw a ball vertically upward a distance of 2.50 m. Given that the acceleration of gravity on the Moon is 1.67 m/s2, how high could you throw a ball on the Moon? (Take the y-axis in the vertical direction, and assume that the location of your hand is at y = 0.)
Answer #1

using 3rd equation of motion

v^2 = 2 gh

v^2 = 2 * 9.8* 2.5

v = 7 m/s

when experiment done on moon

h = v^2/ ( 2 g') = 7^2 / (2 * 1.67)

h = 14.671 m

======

Comment before rate in case any doubt, will reply for sure.. Goodluck